RC Meeting and Run Plan, Monday, January 13th 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Target updates

  • During the previous pumping (before closing up the target enclosure) we observed plateauing of the polarization due to masing effects for the transverse direction. For the longitudinal direction (which had a higher gradient than nominal), we observed similar behavior but with huge AFP losses. This was a common situation encountered by Transversity experiment folks and was addressed in the past by tweaking the magnetic field gradient.
  • Long story short, during the most recent pumping, the longitudinal polarization looked very good (>50+% polarization) but take this number with a grain of salt. Overall, there is some tweaking to do to possibly get higher He3 polarization but nevertheless, we decided to close up the target enclosure and prepare for beam.
  • We are currently taking 1-pass calibration data. We finished density measurements and optics runs. We will continue with our longitudinal and transverse asymmetry measurements for the rest of the swing + owl shift, provided the beam permits us. After that, we will move on to 5-pass tomorrow evening (?).

Remaining work:

  • Update the TO manual and have a walkthrough of the polarimetry software for all the target experts.
  • Announce a TO class for sometime this week.

Current Run Plan

  • Monday: 1-pass data with longitudinal and transverse asymmetries (https://hallcweb.jlab.org/wiki/index.php/1-pass_calibration_for_A1n_(Jan_2020) ). Are we starting off with Transverse or Longitudinal?
  • Tuesday: Finish 1-pass at 8:00 AM, open the target enclosure, remove collimators, do a local EPR and close the panels. Transition to 5-pass.
  • Wednesday: 5-pass beam and Moller measurement
  • Thursday: TO class?

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[1]
  • Temporary Laser instructions can be found here [2]

Ongoing Work in the Hall

Planned Accelerator Configuration Changes

Special Requests

  • Keep target expert schedule filled out on whiteboard